During their physics field trip to the amusement park, Tyler and Maria took a rider on the Whirligig. The Whirligig ride consists of long swings which spin in a circle at relatively high speeds. As part of their lab, Tyler and Maria estimate that the riders travel through a circle with a radius of 6.5 m and make one turn every 5.8 seconds. Determine the speed of the riders on the Whirligig. (Give your answer to the nearest whole number
1
Expert's answer
2020-04-29T09:50:31-0400
"v=\frac{2\pi r}{T}"

"v=\frac{2\pi (6.5)}{5.8}=7.0\frac{m}{s}"